OpenStudy (anonymous):

To which subsets) does the number – √ 25 belong? Choose all that apply. irrational numbers rational numbers integers whole numbers natural numbers

OpenStudy (anonymous):

We know\[ \mathbb{N}\subset\mathbb{N_0}\subset\mathbb{Z}\subset\mathbb{Q}\subset\mathbb{R} \]And, we also know \(-\sqrt{25}=-5\), so, from this, we know, -5 is an integer, which must also be a rational number. But it is not irrational.